OpenStudy (anonymous):

Find the missing terms of the arithmetic sequence 4______ 22 25______,53_______

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Hmm, lets call the missing term x, the difference between a number and the previous one is always the same in an arithmetic sequence. 4,x,22 so x-4 should be the same as 22-x do you follow?

OpenStudy (anonymous):

ehhh not really but some how i got 18

OpenStudy (anonymous):

Thats not right =p say you have the arithmetic sequence 2,4,6,8 8-6=2 6-4=2 4-2=2 so the difference is always the same so in 4,x,22 22-x is the same as x-4 therefore 22-x=x-4 solve for x

OpenStudy (anonymous):

whaai don't get it lol it just more confusing

OpenStudy (anonymous):

how do you solve for x

OpenStudy (anonymous):

Basic algebra add x to both sides 22=2x-4 add 4 to both sides 26=2x divide both sides by 2 13=x

OpenStudy (anonymous):

so 13 would be the answer for the first one and how did you get 2x

OpenStudy (anonymous):

i added x to both sides of the equation x+x is 2x

OpenStudy (anonymous):

ok so is that the answer and do i do the same for the next question

OpenStudy (anonymous):

yes 13 is the answer same idea for the second one, can you try?

OpenStudy (anonymous):

ok and i was gonna try but then its more than two numbers, what now? :O

OpenStudy (anonymous):

just worry about the first blank for now. pretend its 25,x,53

OpenStudy (anonymous):

i got -39 , this is what i did am not sure but 25=2x-53 then divide -78 over 2 which is how i get -39

OpenStudy (anonymous):

you do 53-x=x-25

OpenStudy (anonymous):

omg can you help me with that..?

OpenStudy (anonymous):

Add x to both sides first

OpenStudy (anonymous):

yea and i got 53=2x-25

OpenStudy (anonymous):

thats right, now add 25 to both sides

OpenStudy (anonymous):

i did and overall i got 39

OpenStudy (anonymous):

Thats right, thats the first blank

OpenStudy (anonymous):

yayy ok, so how is the next one done?

OpenStudy (anonymous):

For that one you can just use the difference, you now have 25,39,53 14 is being added each time, 53+14=?

OpenStudy (anonymous):

which would be 67

OpenStudy (anonymous):

Right =D

OpenStudy (anonymous):

yayy thank you..